Sandhill Inc. has seven-year bonds outstanding that pay a 10 percent coupon rate. Investors buying these bonds today can expect to earn a yield to maturity of 6.350 percent. What is the current value of these bonds? Assume annual coupon payments. (Round answer to 2 decimal places, e.g. 15.25.)
Answer #1

Assuming the face value to be $1,000. If it has a different face value, please put it in comments

Coupon = 0.1 * 1000 = 100

Current value = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Current value = 100 * [1 - 1 / (1 + 0.0635)7] / 0.0635 + FV / (1 + 0.0635)7

Current value = 100 * 5.5136 + 649.88654

Current value = $1,201.25